40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 25 10:20:47 2009, Alberto, Update, ABSL, AbsL PLL not able to lock 
    Reply  Wed Nov 25 11:02:54 2009, Alberto, Update, ABSL, AbsL PLL not able to lock 
    Reply  Wed Nov 25 20:14:58 2009, Alberto, Update, ABSL, AbsL PLL not able to lock: problem fixed 
Message ID: 2329     Entry time: Wed Nov 25 11:02:54 2009     In reply to: 2328
Author: Alberto 
Type: Update 
Category: ABSL 
Subject: AbsL PLL not able to lock 

Quote:

Last night something happened on the beat between the PSL beam and the auxiliary NPRO beam, that spoiled the quality of the beating I had before. As a result the PLL has become unable to lock the two lasers.

The amplitude of the beat at the spectrum analyzer has gone down to -40 dBm from -10 that it was earlier. The frequency has also become more unstable so that now it can be seen writhing within tens of KHz.

Meanwhile the power of the single beams at the PLL photodiode hasn't changed, suggesting that the alignment of the two beam didn't change much.

Changes in the efficiency of the beating between the two beams are not unusual. Although that typically affects only the amplitude of the beat and wouldn't explain why also its frequency has become unstable. Tuning the alignment of the PLL optics usually brings the amplitude back, but it was uneffective today.

It looks like something changed in either one of the two beams. In particular the frequency of one of the two lasers has become less stable.

Another strange thing that I've been observing is that the amplitude of the beat goes down (several dBm) as the beat frequency is pushed below 50 MHz. Under 10 MHz it even gets to about -60 dBm.

I noticed the change yesterday evening at about 6pm, while I was taking measurements of the PLL open loop tranfer function and everything was fine. I don't know whether it is just a coincidence or it is somehow related to this, but Jenne and Sanjit had then just rebooted the frame builder.

 I confirm what I said earlier. The amplitude of the beat is -10 dBm at 300MHz. It goes down at lower frequencies. In particular it gets to-60 dBm below 20 MHz. For some strange reason that I couldn't explain the beating efficiency has become poorer at low frequencies.

ELOG V3.1.3-